OpenStudy (anonymous):

what was original price of camera with 26% of usual price and the discount was $96.20 thanks in advance seems everyone is confused as i am

OpenStudy (anonymous):

do you mean that at 26% off it was $96.20?

OpenStudy (anonymous):

the discount was 26% and that discount was $96.20

OpenStudy (anonymous):

how the question is asked is how it is written in the book

OpenStudy (anonymous):

think i will contact my math teacher

OpenStudy (anonymous):

its ok, I figured it out.

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

@mybaby55 do you understand why the difference between "26% of usual price" and "26% discount" is important?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you set it up as a fraction. so \[\frac{ 26% }{ 96.20 } = \frac{ 100% }{ x }\] and then you solve. 26x = 9620 then divide to get x = 370

OpenStudy (whpalmer4):

If the camera sells for 26% of the original price, the price is 0.26*original price If the camera sells with a 26% discount, the price is original price - 0.26*original price

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

it does now thank you

OpenStudy (anonymous):

you're welcome :)

OpenStudy (whpalmer4):

it's more likely, in my opinion, that the camera is selling for a 26% discount. If the discount = $96.20, and the discount is 26% of the original price, original price * 0.26 = $96.20. original price = $96.20/0.26 = $370

OpenStudy (anonymous):

i was doing it with 100% -26% but was not coming up with an answer with next part of equation

OpenStudy (anonymous):

both have same answer just different ways love them both thanks

OpenStudy (whpalmer4):

If the camera sold for 26% of its original price (meaning the discount was 100% - 26% = 74%), then the original price to result in a discount of $96.20 would be original price * 0.74 = $96.20 original price = $96.20/0.74 = $130

OpenStudy (whpalmer4):

So as you can see, very different answers depending on the interpretation of the problem.

OpenStudy (anonymous):

now i can see what you mean

OpenStudy (whpalmer4):

if "26% of usual price" is intended to be "26% OFF usual price" (in other words, a 26% discount), well, this should serve as an illustration of the importance of careful typing/spelling!

OpenStudy (anonymous):

i was just using the 2nd part of the formula when using the %